OpenStudy (anonymous):

can someone help me find the maximum value for f(x) = -x^2 - 2 x + 24. here is what I found so far: the parabola has a y intercept at (0,24) and 2 intercepts that occur at -6 and 4

OpenStudy (anonymous):

You want to find the vertex. Do you know how to do this?

OpenStudy (amistre64):

the maximum value of an upside down parabola; is the vertex

OpenStudy (amistre64):

or the derivative depending on what class your in :)

OpenStudy (anonymous):

the vertex cordinates are ( -1, and ?)

OpenStudy (amistre64):

plug in x=-1 and solve for y

OpenStudy (anonymous):

i didnt find the other one

OpenStudy (anonymous):

plug into the original function?

OpenStudy (amistre64):

yes: y= (.....)

OpenStudy (amistre64):

what does y= when x=-1? well, we use the equation to determine that :)

OpenStudy (anonymous):

Yes the vertex is at: \[\huge (\frac{-b}{2a},f(\frac{-b}{2a}))\] when the function is written as ax^2+bx+c

OpenStudy (amistre64):

or we can maipulat eit into a vertex form ...

OpenStudy (amistre64):

manipulate even ....

OpenStudy (anonymous):

yes, but you'd have to complete the square. easier in its current form :)

OpenStudy (anonymous):

well this one is actually pretty easy

OpenStudy (amistre64):

f(x) = -x^2 - 2 x + 24 f(x) = -(x^2 + 2 x +1) -1 + 24 f(x) = -(x+1)^2 +23 right?

OpenStudy (anonymous):

ok i found y=27 when x=-1

OpenStudy (amistre64):

lets chk that then :) f(-1) = -(-1)^2 - 2(-1) + 24 = -1 +2 + 24 = 25 might be better

OpenStudy (anonymous):

ok so ( -1,25) are my vertex

OpenStudy (amistre64):

i pulled out a -1 in my vertex form instead of the proper +1 .... f(x) = -(x+1)^2 + 25 would have been gooer

OpenStudy (amistre64):

yes, and the vertex tells us the "highest" point in this case

OpenStudy (anonymous):

so the highest point is the min and the lowest point is the max

OpenStudy (amistre64):

let me draw it

OpenStudy (amistre64):



OpenStudy (amistre64):

the vertex point is our highest point on this f(x) so the highest value we get is: 25

OpenStudy (anonymous):

Yes, when the coefficient of x^2 is negative, the parabola opens down and has a maximum at the vertex; when it's positive, the parabola opens up and has a minimum at the vertex

OpenStudy (anonymous):

so the maxis 25?

OpenStudy (anonymous):

yes:25

OpenStudy (amistre64):

unless we found a different meaing for maximum and 25 ... id would say that yes; the maximum is 25

OpenStudy (anonymous):

so find the vertex's then plug it back into the function to find the minimum and maximum?

OpenStudy (amistre64):

what would minimum suggest to you? smallest possible value right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

the minimum of this function is then: -infinity

OpenStudy (anonymous):

yes so it would be concave up

OpenStudy (amistre64):

cave down

OpenStudy (amistre64):